A 57.1 g piece of zinc (Csp = 0.388 J/g · °C), initially at 285.8 °C, is added to 155.3 g of a liquid, initially at 22.1 °C, in an insulated container. The final temperature of the metal–liquid mixture at equilibrium is 36.7 °C. What is the identity of the liquid? Neglect the heat capacity of the container.
Answer #1

Liquid = unknown metal = zinc Cp = 0.3883/9.c mass = 57-19 Ti = 285.8c Cp = ? mass = 155.39 : 21) Ti= 22-ico If = 36.7°C If =